OpenStudy (anonymous):

What is the relative extrema for the function f(x) = 1/4 (x^4) -8x^2?

myininaya (myininaya):

Find f'

myininaya (myininaya):

set f'=0 and find where f' dne exist and use these values as critical numbers since that is the way they are defined.

OpenStudy (anonymous):

what is F^1 of the problem?

OpenStudy (anonymous):

Finf f'(x) and determine where it is zero or undefined. Then use First Derivative Test. f ' (x) =x^3 - 16x = x(x^2 - 16) = x(x-4)(x+4) So f ' (x) = 0 when x = 0, 4, and -4 Now use the First Derivative Test.

myininaya (myininaya):

You can use the first derivative test to find the relative extrema. Or you can use the second derivative. f prime

OpenStudy (anonymous):

what is the first derivate test and how would I apply it?

OpenStudy (anonymous):

It says that where f' switches from negative to positive, there will exist a rel min; from positive to negative, will reult in a rel max. You want to test the intervals: x< - 4 -4 < x < 0 0 < x < 4 x > 4 Check these 4 intervals and see if the derivative is positive or negative in that interval. Then see if it switches from + to -, or - to +. If it switches from + to + or - to -, then no rel max/min exists there.

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

Got it. thank you